### v4

# LatentExplainer: Explaining Latent Representations in Deep Generative Models with Multi-modal Foundation Models

Abstract
Deep generative models like VAEs and diffusion models have advanced various generation tasks by leveraging latent variables to learn data distributions and generate high-quality samples. Despite the field of explainable AI making strides in interpreting machine learning models, understanding latent variables in generative models remains challenging. This paper introduces LatentExplainer, a framework for automatically generating semantically meaningful explanations of latent variables in deep generative models. LatentExplainer tackles three main challenges: inferring the meaning of latent variables, aligning explanations with inductive biases, and handling varying degrees of explainability. Our approach perturbs latent variables, interpreting changes in generated data, and uses multi-modal large language models (MLLMs) to produce human-understandable explanations. We evaluate our proposed method on several real-world and synthetic datasets, and the results demonstrate superior performance in generating high-quality explanations for latent variables. The results highlight the effectiveness of incorporating inductive biases and uncertainty quantification, significantly enhancing model interpretability.
1 Introduction
Deep generative models, such as Variational Autoencoders (VAEs) Kingma and Welling (2013) and diffusion models Rombach et al. (2022), have become a state-of-the-art approach in various generation tasks Razavi et al. (2019); Ho et al. (2022); Yang et al. (2023). This method effectively leverages latent variables to learn underlying data distributions and generate high-quality samples by capturing the underlying structure of high-dimensional data in a low-dimensional semantic space Vahdat et al. (2021). To enhance performance, inductive biases are often enforced over latent variables. For instance, disentanglement is a rule of thumb which enforces orthogonality among different latent variables Ding et al. (2020). Moreover, sometimes latent variables can be grouped, leading to combination bias Klys et al. (2018); Guo et al. (2020). More recently, the desire for controllability in deep generative models, where latent variables are associated with specific properties of interest, Wang et al. (2024), has given rise to conditional bias. As the latent variables represent all the information in a lower dimension, they can be considered as an effective abstraction of key factors in the data. Therefore, it is critical to develop methods for automatically decomposing and explaining meaningful latent dimension semantics given a pretrained generative model and its inherent inductive biases, as illustrated in Figure 1. Such capabilities are essential for applications in controllable data generation, interpretable model explanations, and natural language understanding.
Figure 1: Illustration of how a pretrained generative model, guided by inductive bias formulas, automatically decodes and interprets meaningful latent dimension semantics. Zoom in for details.
The field of explainable artificial intelligence (XAI) has extensively investigated the interpretation of machine learning models Adadi and Berrada (2018). However, interpreting latent variables in deep generative models remains underexplored. Machine learning model explanation, a.k.a., post-hoc explanation, can be categorized into global and local explanations Gao et al. (2024). Global explanations focus on elucidating the entire model, while local explanations target the reasoning behind specific predictions. Global explanations are more challenging, with existing work mostly emphasizing attributions to identify which features are most important for model decision-making Saleem et al. (2022). The missing piece is understanding the meaning of features when they are unknown, which is very common in deep generative models. More recently, one category of global explanation methods, called concept-based explanations, aims to generate more human-understandable concepts as explanations Poeta et al. (2023). However, current concept-based methods often rely on human heuristics or predefined concept and feature space, limiting the expressiveness of the explanations and falling short of achieving truly automatic explanation generation Kim et al. (2018); Koh et al. (2020); Bai et al. (2022).
Despite the progress in XAI, interpreting latent variables in deep generative models presents significant challenges. First, these variables are not grounded in real-world concepts, and the black-box nature of the models prevents us from inferring the meaning of latent variables from observations. Second, explanations must adhere to the inductive biases imposed on the latent variables, which is essential yet difficult to ensure. For example, in disentangled latent variables, the semantic meanings should be orthogonal. Third, different latent variables have varying degrees of explainability. Some may be trivial to data generation and intrinsically lack semantic meaning. It is crucial to identify which latent variables are explainable and which do not need explanations.
To address the aforementioned challenges, we propose LatentExplainer, a novel and generic framework that automatically generates semantically-meaningful explanations of latent variables in deep generative models. Specifically, to explain these variables and work around the black-box nature of the models (Solution to Challenge 1), we propose to perturb each latent variable and explain the resulting changes in the generated data. Specifically, we perturb and decode each manipulated latent variable to produce the corresponding sequence of generated data samples. The trend in the sequence is leveraged to reflect the semantics of the latent variable to be explained. To align explanations with the intrinsic nature of the deep generative models (Solution to Challenge 2), we design a generic framework that formulates inductive biases on the Bayesian network of latent variable models into textual prompts. These prompts are understandable to large foundation models and humans. To handle the varying degree of explainability in latent variables (Solution to Challenge 3), we propose to measure the confidence of the explanations by estimating their uncertainty. This approach assesses whether the latent representations are interpretable and selects the most consistent explanations, ensuring accurate and meaningful interpretation of the latent variables.
2 Related Work
Deep Generative Models. Deep generative models are essential for modeling complex data distributions. Variational Autoencoders (VAEs) are prominent in this area, introduced by Kingma and Welling Kingma and Welling (2013). VAEs encode input data into a latent space and decode it back, optimizing a balance between reconstruction error and the Kullback-Leibler divergence Rezende et al. (2014). They have diverse applications, including image generation Yan et al. (2016), semi-supervised learning Kingma et al. (2014), and anomaly detection An and Cho (2015).
Diffusion models, proposed by Ho et al. Ho et al. (2020), generate data by a diffusion process that gradually adds noise to the data and then learns to reverse this process to recover the original data. These models have achieved high-fidelity image generation, surpassing GANs in quality and diversity. Latent diffusion models allow the model to operate in a lower-dimensional space, which significantly reduces computational requirements while maintaining the quality of the generated samples Rombach et al. (2022). Advances have made them applicable to text-to-image synthesis Nichol and Dhariwal (2021), audio generation Kong et al. (2020), and molecular design Xu et al. (2022).
Latent Variable Manipulation and Explanations. Manipulating latent variables in generative models like VAEs and diffusion models is an important technique for editing and enhancing generated images. A key method is latent traverse, which involves traversing different values of latent variables to achieve diverse manipulations in the generated outputs. This technique allows for precise control over the attributes in generated images, enabling adjustments Chen et al. (2016); Higgins et al. (2017). For example, latent traverse has been effectively employed to disentangle and control various attributes in generated images Brock et al. (2016); Zhu et al. (2016). However, latent traverse is often used for visualization and editing purposes. It has not yet been widely explored as a tool for explaining the underlying latent space. Their explanations primarily rely on using predefined training attributes as text labels or manually adding explanations Shen et al. (2020); Esser et al. (2020). Some concept-based models control latent variables in generative models using category concepts to generate data Tran et al. (2022); Bouchacourt et al. (2018). However, these approaches cannot automatically generate free-form textual explanations.
More recently, Multi-modal Large Language Models (MLLMs) integrate diverse data modalities, enhancing their ability to understand and generate complex information Yin et al. (2023); Bai et al. (2024). Notable models include GPT-4o, which extends GPT-4v with better visual capabilities OpenAI (2024), and Gemini that is a family of highly capable multimodal models Team et al. (2023). We envision leveraging MLLMs to automatically generate explanations for latent variables. Our work focuses on: (1) how to develop prompts that are aligned with the underlying generative model and can be easily understood by MLLMs, as well as (2) how to evaluate the quality of the generated explanations for latent variables.
Figure 2: The overview of our proposed framework LatentExplainer (a) Inductive-bias-guided Data Manipulation generates image sequences by manipulating latent variables with predefined biases; (b) Automatic Prompt Generation with symbol-to-word mapping uses these images and formulas to create prompts for a MLLM to produce explanations; (c) Uncertainty-aware Explanation Generation evaluates multiple responses from the MLLM, selecting the most consistent explanation with a certainty score.
3 Preliminaries and Problem Formulation
Deep Generative Models. These are a class of models that learn a mapping between observations $\mathbf{x}$ and key underlying factors $\mathbf{z}$. These models are widely used in deep generative models such as VAEs and latent diffusion models. VAEs, for instance, introduce a probabilistic approach to encoding data by maximizing the evidence lower bound (ELBO) Kingma and Welling (2013):
$\mathcal{L}_{\text{ELBO}}=\mathbb{E}_{q(\mathbf{z}|\mathbf{x})}[\log p(\mathbf{x}|\mathbf{z})]-\text{KL}(q(\mathbf{z}|\mathbf{x})\|p(\mathbf{z})),\vspace{-2mm}$
where KL stands for Kullback–Leibler divergence. Latent diffusion models further refine the generation process by iteratively refining noise into structured data Rombach et al. (2022). These models effectively capture the underlying structure of data in a low-dimensional semantic space.
Latent variable manipulation of diffusion models aims at transversing the latent representation $\mathbf{z}$ along the semantic latent direction $\mathbf{v}$. The perturbed vector $\tilde{\mathbf{z}}=\mathbf{z}+\gamma\left[\mathcal{G}(\mathbf{z}+\mathbf{v})-\mathcal{G}(\mathbf{z})\right]$, where $\gamma$ is a hyper-parameter controlling the strength Park et al. (2023). An image sequence $\mathcal{S}$ can then be generated by $\mathcal{G}(\mathbf{\tilde{z}})$. The perturbations in the semantic latent direction lead to semantic changes in the generated image sequence.
Inductive Bias in Latent Variables. Inductive biases are commonly imposed on latent variables to enhance the performance and interpretability of deep generative models. These biases can be categorized into several types: Disentanglement Bias: It enforces orthogonality among different latent variables, ensuring that each latent variable captures a distinct factor of variation in the data Ding et al. (2020). Combination bias: Sometimes latent variables are grouped, leading to biases in how they interact and combine to represent complex data structures Klys et al. (2018). Conditional Bias: It emphasizes the relationship between specific properties of interest and the corresponding latent variables Guo et al. (2020). Other Biases: Various other biases can be introduced depending on the specific requirements of the model to guide the learning process more effectively.
Problem Formulation. We assume a dataset $\mathcal{D}$, where each sample consists of $x$ or $(x,y)$, with $x\in\mathbb{R}^{N}$ and $y=\left\{y_{k}\in\right.$$\mathbb{R}\}_{k=1}^{K}$ as $K$ properties of $x$. The dataset $\mathcal{D}$ is generated by M latent variables $z_{i}$, where $i\in\{1,\ldots,M\}$. $z_{i}$ can be a single latent variable or a group of correlated latent variables. Suppose we are given a generative model with a set of formulas $\mathcal{F}$ with respect to $z_{i}$, where $\mathcal{F}$ represents an inductive bias that the generative model must satisfy. Our goal is to derive a textual sequence that explains the semantic meanings of the latent variable $z_{i}$.
4 Proposed Method
4.1 Overview of LatentExplainer
This paper focuses on the tasks in explaining the semantics of latent variables $\{{z}_{i}\}_{i=1}^{M}$ in deep generative models. To interpret the semantics of latent variables and work around the blackbox nature of deep generative models, we propose to perturb each latent variable and explain the change it imposes on the generated data. To solve these challenges, we propose a novel LatentExplainer scheme. We perturb $z_{i}$ and subsequently decode the manipulated latent variables into generated data that are perceptible by humans such as an image. Through a series of perturbations on $z_{i}$, a sequence of generated data samples can be obtained to reflect the changes in $z_{i}$ in Figure 2(a). When explaining latent variable models, it is crucial to fully leverage and align with the prior knowledge about them. To do this, we design a generic framework that can automatically formulate inductive bias of generative models into textual prompts. Specifically, we have summarized three common inductive biases and designed their symbol-to-word one-on-one prompts $\mathcal{P}$ (Section 4.2). Our scheme can adaptively convert any user-provided inductive bias formulas $\mathcal{F}$ into a corresponding prompt $\mathcal{P}$ to provide more accurate explanations of the latent representations in Figure 2(b) (Section 4.3). Eventually, the explanations are selected through an uncertainty quantification approach to assess whether the latent representations are interpretable and select the most consistent explanations in Figure 2(c) (Section 4.4).
4.2 Inductive-bias-guided Prompt Framework
4.2.1 Generic Framework
In this section, we propose a generic framework that can verbalize the inductive bias in deep generative models into prompts for better latent variable explanations. The prevalent inductive biases in deep generative models are categorized into three types: disentanglement bias, conditional bias, and combination bias. Intrinsically, they are mathematical expressions of the conditional independence among the latent variables, which can confine their explanations but cannot be directly input into large foundation models.
Framework: Our framework proposes a principled, automatic way that translate the mathematical expression to textual prompts. The prompts include adaptive prompts and a fixed ending. The adaptive prompts are converted from the inductive bias formulas. The formulas contain mathematical symbols that consist of mathematical variables and mathematical operators. We use the same color to represent the correspondence between mathematical symbols in the formulas and the text in the prompts. The translation mechanism of adaptive prompts is shown in Table 1.
Grammar # Symbol Prompt
1 $p$($z_{i}$$\cdot$ ⋅) pattern of change
2 $p(z_{i}\mid$$z_{i^{\prime}}$), $\forall i\neq i^{\prime}$ other variations
3 $p_{k}$ property of interests
4 $G$ a group
5 $\in$ associated with
6 $\notin$ not associated with
7 = same
8 $\neq$ change
Table 1: Lookup table for symbol-to-word mapping.
Fixed Ending: What is the pattern of change? Write in a sentence. If there is no clear pattern, just write "No clear explanation".
4.2.2 From Disentanglement Bias to Prompts
Disentanglement bias refers to the model’s ability to separate independent factors in the data Ding et al. (2020); Wu et al. (2023). The formula representing this bias focuses on ensuring that different latent variables correspond to different independent underlying factors. Independent factors would be invariant with respect to one another (Ridgeway, 2016). By disentangling these factors, researchers can better understand the underlying structure of the data and improve the model’s performance on tasks such as representation learning.
Formula:
$p({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}z_{i^{\prime}}=\alpha}){\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}=}p({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}z_{i^{\prime}}=\beta}),{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\forall i\neq i^{\prime},\alpha\neq\beta}.\vspace{-1mm}$
The above formula is translated into the following prompting using the grammar #1,2,7.
Prompt: These two rows of images show the same pattern of change despite other variations.
4.2.3 From Combination Bias to Prompts
Combination bias involves understanding how different latent variables interact within groups and remain independent across groups Klys et al. (2018). This bias is significant as it helps in identifying how combinations of factors contribute to the overall data generation process. Recognizing these interactions enables researchers to design models that can generate more complex and realistic data by capturing intricate relationships within the data.
  • No inter-group correlation:
    Formula: $p\left({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}z_{j}=\alpha}\right){\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}=}p\left({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}z_{j}=\beta}\right),$
    $\forall{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\in}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}G},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}z_{j}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\in}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}G^{\prime}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}G\neq G^{\prime}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\alpha\neq\beta}.$
The above formula is translated into the following prompting using the grammar #1,2,4,5,7. Prompt: The pattern of change is associated with a group. The first two rows of images show the same pattern of change despite other variations in another group.
  • Intra-group correlation:
    $\begin{split}\noindent\textbf{Formula:}\quad&p\left({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}z_{i^{\prime}}=\alpha}\right){\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\neq}p\left({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}z_{i^{\prime}}=\beta}\right),\\ &\forall{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}z_{i}^{\prime}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\in}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}G},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}i\neq i^{\prime},\alpha\neq\beta}.\end{split}$
The above formula is translated into the following prompting using the grammar #1,2,4,5,8.
Prompt: The pattern of change is associated with a group. The pattern of change in the last two rows of images should change given other variations.
4.2.4 From Conditional Bias to Prompts
Conditional bias focuses on the relationship between specific properties of interest and the corresponding latent variables. This bias is important because it allows models to generate data conditioned on particular attributes, enhancing the model’s ability to produce targeted and controlled outputs.
Formula:
$p({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}p_{k}=\alpha}){\color[rgb]{0,1,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,1}\pgfsys@color@cmyk@stroke{1}{0}{0}{0}\pgfsys@color@cmyk@fill{1}{0}{0}{0}\neq}p({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}p_{k}=\beta}),\forall{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{i}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\in}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}G}{\color[rgb]{1,.5,0}{}_{k}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\alpha\neq\beta}.\vspace{-2mm}$
The above formula is translated into the following prompting using the grammar #1,2,3,4,5,8.
Prompt: If the pattern of change is associated with the group of the property of interest, this image sequence will change as other variations in $[\text{property}_{k}]$.
There may exists a latent variable $z_{j}$ that are independent of $p_{k}$:
$p({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{j}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}p_{k}=\alpha}){\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}=}p({\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{j}}\mid{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}p_{k}=\beta}),\forall{\color[rgb]{.75,0,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,0,.25}z_{j}}{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\notin}{\color[rgb]{1,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,1}\pgfsys@color@cmyk@stroke{0}{1}{0}{0}\pgfsys@color@cmyk@fill{0}{1}{0}{0}G}{\color[rgb]{1,.5,0}{}_{k}},{\color[rgb]{.75,.5,.25}\definecolor[named]{pgfstrokecolor}{rgb}{.75,.5,.25}\alpha\neq\beta}.\vspace{-2mm}$
The above formula is translated into the following prompting using the grammar #1,2,3,4,6,7.
Prompt: If the pattern of change is not associated with the group of the property of interest, this image sequence will remain constant despite other variations in $[\text{property}_{k}]$.
Algorithm 1 Generate Prompt from Inductive Bias by LatentExplainer
1:Inductive Bias Formula(s) $\mathcal{F}$, Few-Shot Examples $\mathcal{H}$, pre-trained LLM $\pi_{\theta}$, an optional input of the information about the symbol $\mathcal{I}$.
2:Generated Prompt $\mathcal{P}$
3:function $\mathcal{F}$, $\mathcal{H}$, $\pi_{\theta}$)
4:     symbols $\leftarrow$ $\mathcal{F}$)
[…]
6:     for symbol in symbols do
7:         semantic $\leftarrow$ $\pi_{\theta}([{\text{symbol},\mathcal{H},\mathcal{I}}])$
8:         semantics[symbol] $\leftarrow$ semantic
9:     end for
10:     $\mathcal{P}=\pi_{\theta}([\mathcal{F},\mathcal{H},\text{semantics}])$
11:     return $\mathcal{P}$
12:end function
4.3 Automatic In-context Prompt Generation
By leveraging these three common inductive biases identified in generative models, we can automatically generate prompts $\mathcal{P}$ that align with $\mathcal{F}$ using in-context learning. The pseudo-code of this section can be found at Algorithm 1 and a more detailed example is at Appendix  C.
Our approach starts by extracting mathematical symbols from the given formula $\mathcal{F}$ using the ExtractSymbols function (line 2). This function traverses the formula to identify and isolate the mathematical symbols.
Next, the algorithm initializes an empty dictionary semantics to store the semantic representations of these symbols (line 3). For each symbol, the pre-trained LLM $\pi_{\theta}$ extracts its semantic meaning based on few-shot examples $\mathcal{H}$ and the optional input symbol information $\mathcal{I}$ (lines 4-6).
Finally, the algorithm generates the prompt $\mathcal{P}$ using the formula $\mathcal{F}$, the few-shot examples $\mathcal{H}$, and the gathered semantics (line 8). This step-by-step reasoning process ensures that the generated prompts are contextually relevant and aligned with the underlying formulas, which could reduce hallucination and enhance model performance.
4.4 Uncertainty-aware Explanations
Uncertainty-aware methods can be applied to large language model responses Lin et al. (2023) and image explanations Zhao et al. (2024). To measure the uncertainty of the responses from GPT-4o, we sampled $n$ times from the GPT-4o to generate the responses =$\mathcal{R}=\{r_{1},r_{2},r_{3},...,r_{n}\}.$ The certainty score of the explanation is the average pairwise cosine similarity of the responses $\mathcal{R}$: $1/C\cdot{\sum}_{i=1}^{n}{\sum}_{j=1,i\neq j}^{n}\text{sim}(r_{i},r_{j})$ where $C=n\cdot(n-1)$. The implementation details is in Appendix B.
5 Experiment
The experiments are conducted on a 64-bit machine with 24-core Intel 13th Gen Core i9-13900K @ 5.80GHz, 32GB memory and NVIDIA GeForce RTX 4090. We use GPT-4o as our MLLM and set $temperature=1$ and $top\_p=1.$ The code is available at  https://anonymous.4open.science/r/LatentExplainer-89AF.
5.1 Dataset
We utilized five datasets to evaluate the performance of different generative models under three inductive biases: CelebA-HQ (Karras et al., 2018), AFHQ (Choi et al., 2020), LSUN-Church (Yu et al., 2015) for the unconditional and conditional diffusion models, and CelebA-HQ (Karras et al., 2018), 3DShapes (Burgess and Kim, 2018), and dSprites (Matthey et al., 2017) for the VAE models. The CelebA-HQ dataset is a high-quality version of the CelebA dataset, consisting of 30K images of celebrities, divided into 28K for training and 2K for testing. The LSUN-Church dataset contains large-scale images of church buildings. The AFHQ dataset includes high-quality images of animals, divided into three categories: cats, dogs, and wild animals. 3DShapes is a synthetic dataset contains images of 3D shapes with six factors of variation: floor hue, wall hue, object hue, object scale, object shape, and wall orientation, divided into 384K for training and 96K for testing. dSprites consists of 2D shapes (hearts, squares, ellipses) generated with five factors of variation: shape, scale, orientation, position X, and position Y, divided into 516K for training and 221K for testing.
Model Method BLEU@4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BERTScore $\uparrow$ BARTScore $\uparrow$
DDPM GPT-4o Baseline 2.90 13.81 11.84 7.08 87.33 -3.32
+ LatentExplainer w/o IB 3.70 14.22 12.17 7.90 87.30 -3.31
+ LatentExplainer w/o UQ 16.81 31.99 29.04 19.79 89.88 -3.08
+ LatentExplainer 25.11 39.11 36.89 25.97 91.19 -2.98
$\beta$-TCVAE GPT-4o Baseline 5.53 27.70 19.43 10.73 88.97 -3.14
+ LatentExplainer w/o IB 3.85 27.71 20.68 10.55 88.93 -3.15
+ LatentExplainer w/o UQ 18.70 41.13 33.84 22.63 91.03 -2.86
+ LatentExplainer 22.63 47.35 35.63 24.88 91.91 -2.80
Table 2: Quantitative explanation comparison of the disentanglement bias.
Model Method BLEU@4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BERTScore $\uparrow$ BARTScore $\uparrow$
CSVAE GPT-4o Baseline 5.79 28.03 20.24 10.58 89.05 -3.02
+ LatentExplainer w/o IB 8.73 27.32 21.25 12.18 89.16 -3.03
+ LatentExplainer w/o UQ 16.21 35.60 27.58 22.81 90.57 -2.80
+ LatentExplainer 25.85 41.54 35.22 28.51 91.29 -2.74
Table 3: Quantitative explanation comparison of the combination bias.
Model Method BLEU@4 $\uparrow$ ROUGE-L $\uparrow$ METEOR $\uparrow$ SPICE $\uparrow$ BERTScore $\uparrow$ BARTScore $\uparrow$
Stable Diffusion GPT-4o Baseline 7.07 25.65 21.79 14.72 89.70 -3.27
+ LatentExplainer w/o IB 8.35 26.32 22.94 14.77 89.61 -3.25
+ LatentExplainer w/o UQ 14.85 32.83 27.03 22.02 90.61 -3.12
+ LatentExplainer 16.70 36.32 29.97 23.00 90.97 -3.08
CSVAE GPT-4o Baseline 6.48 22.98 17.03 8.55 88.49 -3.11
+ LatentExplainer w/o IB 9.21 26.68 21.09 11.04 88.98 -3.05
+ LatentExplainer w/o UQ 11.82 29.56 22.07 14.20 88.89 -3.01
+ LatentExplainer 23.70 39.29 32.16 20.33 90.53 -2.89
Table 4: Quantitative explanation comparison of the conditional bias.
5.2 Generative Models under Inductive Biases
We explore the latent space in generative models that satisfy the aforementioned three types of inductive biases. For each type, we present the relevant generative models that align with the corresponding inductive bias: (1) Disentanglement Bias: $\beta$-TCVAE (Chen et al., 2018) explicitly penalizes the total correlation of the latent variables to disentangle the latent representations. Denoising Diffusion Probabilistic Model (DDPM) (Ho et al., 2020) adds Gaussian noise independently at each timestep in the forward process and eventually transforms into pure Gaussian noise, in which the covariance matrix is diagonal. This assumes the latent factors are independent; (2) Combination Bias: CSVAE (Klys et al., 2018) has two groups of latent variables $z$ and $w$, where $z$ and $w$ are uncorrelated and the latent variables within the group are correlated; (3) Conditional Bias: CSVAE also satisfies conditional bias because one group of latent variables $w$ is associated with the properties while the other group of latent variables $z$ minimizes the mutual information with the properties. Stable Diffusion (Rombach et al., 2022) is a latent diffusion model to generate images conditioned on prompts.
5.3 Human Annotaions
The ground-truth annotations of explanations are performed by four annotators from the United States and China. All annotators are students that had at least an undergraduate degree. Annotators were presented with the same images and the prompts as MLLM and were asked to annotate the pattern of the images. If there is no clear pattern, just write "No clear explanation". The annotations are then aggregated as references to calculate the automated evaluation metrics.
5.4 Quantitative Analysis
For the quantitative explanation evaluation, we use BLEU Papineni et al. (2002), ROUGE-L Lin (2004), METEOR Banerjee and Lavie (2005), SPICE Anderson et al. (2016), BERTScoreZhang et al. (2019), and BARTScore Yuan et al. (2021) as the automated metrics to assess the generated explanations. BLEU, ROUGE-L, METEOR are n-gram-based metrics that measure the overlap between generated and reference texts. SPICE compares scene graphs derived from the generated and reference texts. BERTScore and BARTScore utilize pre-trained transformer-based language models to compute contextual embeddings of the generated and reference texts. These metrics together provide a comprehensive assessment of both the lexical and semantic quality of the explanations.
To understand the impact of various components in our proposed LatentExplainer, we conducted a comprehensive quantitative analysis across different models (DDPM, $\beta$-TCVAE, Stable Diffusion, and CSVAE) and inductive biases (disentanglement, combination, and conditional). Specifically, we compare the removal of inductive bias prompts (IB), the removal of uncertainty quantification (UQ), and the full model against the baseline GPT-4o. The results are weighted means across three datasets in Tables 2, Tables 3, and Tables 4. The specific results for each dataset are provided in Table 5 and Table 6 in the Appendix. Removing inductive bias prompts leads to a substantial drop in all generative models. Their consistent results demonstrate that inductive bias is the most important and necessary component when explaining the latent representations of generative models. The removal of uncertainty quantification also results in a slightly decreased performance in all generative models, indicating that uncertainty quantification is also effective, though not as critical as inductive bias prompts. The full model, which incorporates both inductive bias prompts and uncertainty quantification, achieves the highest overall performance, and outperforms all baselines across all models. This confirms the necessity of both inductive bias prompts and uncertainty quantification in our LatentExplainer framework, demonstrating their significant contributions to improving explanatory performance across various generative models. It also shows that our framework can effectively design prompts for different inductive biases in generative models to improve the accuracy of latent explanations.
Figure 3: Visualization of the generated explanations with the inductive bias prompt for the disentanglement bias.
Figure 4: Visualization of the generated explanations w/o the inductive bias prompt for the disentanglement bias.
5.5 Qualitative Evaluation
To analyze the explanations for DDPM under disentanglement bias, we manipulate the latent representation along a latent direction and compare it with the one that first traverses along another latent direction and then traverses along the same latent direction. The disentangled latent variable would be invariant with respect to the variations in another latent dimension. In Figure 3, for each latent direction, we pass two image sequences and an inductive bias prompt based on the disentanglement formula to the MLLM to obtain a common latent explanation. In comparison, the explanations generated without the inductive bias prompt as shown in Figure 4 tend to show "no clear explanation" or wrong explanations as they only align with one image sequence but do not reflect the common pattern in both image sequences. The addition of the inductive bias prompts can assist with ruling out the variation effects of other latent variables to capture the actual meaning.
We also qualitatively evaluate the explanations for CSVAE under combination bias. We transverse a latent variable and compare it with the one that first traverses another latent variable in another group and then traverses the same latent variable, which is similar to the disentanglement bias. We then compare with the one that first transverses another latent variable in the same group and then traverses the same latent variable. As Figure 10 depicts, our model can clearly show the explanations of latent variable $z_{1},z_{2},z_{3}$ as the color of the ground, the background color, and the shape of the object. The effect of removing inductive bias prompts leads to no clear explanation or wrong explanations like the disentanglement bias.
In Figure 11, we provide the "young appearance" prompt to Stable Diffusion under conditional bias, and the explanations of all three top latent directions reflect the meaning of youth. The addition of inductive bias prompts can better identify the relation with the property of interest to capture the actual meaning of latent variables. In comparison, the one without the inductive bias prompts in Figure 12, cannot find clear explanations or simply describe the characteristics in the image sequence, lacking an abstract generalization. More qualitative evaluation results can be found in Appendix D.
6 Conclusion
In this paper, we introduced LatentExplainer, a framework designed to generate semantically meaningful explanations of latent variables in deep generative models. Our experiments quantitatively revealed that the inclusion of inductive bias prompts significantly enhances the quality of explanations generated by LatentExplainer. In qualitative evaluations, LatentExplainer provided clear explanations for the latent variables, effectively reflecting superior performance when inductive bias prompts were utilized. The quantitative analysis further confirmed the necessity of inductive bias prompts and uncertainty quantification in producing accurate and consistent explanations.
7 Limitation
In this work, we use GPT-4o as our MLLM agent to explain the latent representations. Although GPT-4o has improved vision capabilities across most tasks, we still find it incapable of identifying intricate facial features, positions, and orientations. It is necessary to enhance visual understanding to improve the explanation quality for MLLM.
Appendix A Latent Variable Perturbation
For VAE models, every time we target at each latent variable. The targeted latent variable is changed from -3 to 3, which is the 3 times the standard deviation of the standard Gaussian distribution of a latent variable, with a step size of 1.5. All the other latent variables are unchanged. For diffusion models, we follow $\tilde{\mathbf{z}}=\mathbf{z}+\gamma\left[\mathcal{G}(\mathbf{z}+\mathbf{v})-\mathcal{G}(\mathbf{z})\right]$, where $\gamma$ is the strength. We set the value of $\gamma$ to 0.1, 0.2, 0.3, 0.4, and 0.5 for DDPM, and to 1, 2, 3, 4, and 5 for Stable Diffusion.
Appendix B Uncertainty-aware Implementations
We denote the true label of the interpretability for the i-th latent variable $y_{i}$ as 1 if at least two of the three annotators can see a clear pattern in the generated images, otherwise we denote it as 0. The certainty score of the explanation for the i-th latent variable is $\bar{s}_{i}$. We adopt the Jaccard Index to measure similarity between the predicted label and the true label,
$\begin{split}&E(\varepsilon)=\texttt{Jaccard}(f(\bar{s}_{i},\varepsilon),y_{i}),\text{ where}\\ &f(\bar{s}_{i},\varepsilon)=\mathbbm{1}(\bar{s}_{i}\geq\varepsilon).\end{split}$
The threshold is then selected as the one with the maximum similarity with the true label,
$\varepsilon^{*}=\arg\max_{\varepsilon}E(\varepsilon).$
By solving this equation across all datasets, we find the threshold $\varepsilon=0.2617$. Since $\varepsilon$ is solved across all datasets, it reduces the dependence on any specific dataset and can be applied to new data. Our final output $\mathcal{T}$ is the response that has the highest mean pairwise cosine similarity with other responses if the certainty score is equal or greater than the threshold $\varepsilon$. Otherwise, $\mathcal{T}$ will be "no clear explanation".
Appendix C An example of automatically generating prompts from an inductive bias formula
Example: Given a formula $\mathcal{F}$ which is “$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right),\forall t_{1}\neq t_{2}"$, generate a prompt that reflects the inductive bias.
Solution:
Step 1: Extracting mathematical symbols: Extracting all mathematical symbols from the given formula “$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right),\forall t_{1}\neq t_{2}"$. Symbols extracted are $\mathcal{S}=\left[\right.$$p\left(z_{i}\mid\cdot\right)"$, “$\left.t=t_{1},t=t_{2},\forall t_{1}\neq t_{2}",``\neq"\right]$.
[…]
Step 3: Generating a prompt for the given formula $\mathcal{F}$ : We can generate a prompt for the given formula $\mathcal{F}$ using a pretrained LLM $\pi_{\theta}$ with the following prompt that consists of the few-shot examples $\mathcal{H}$ of the three inductive biases, $\mathcal{F}$, and the gathered $\mathtt{semantics}$:
Formula:
$p\left(z_{i}\mid z_{i^{\prime}}=\alpha\right)=p\left(z_{i}\mid z_{i^{\prime}}=\beta\right),\forall i\neq i^{\prime},\alpha\neq\beta$.
Prompt: These two rows of images show the same pattern of change despite other variations. What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation".
Formula:
$p\left(z_{i}\mid z_{j}=\alpha\right)=p\left(z_{i}\mid z_{j}=\beta\right),\forall z_{i}\in G,z_{j}\in G^{\prime},G\neq G^{\prime},\alpha\neq\beta$.
$p\left(z_{i}\mid z_{i^{\prime}}=\alpha\right)\neq p\left(z_{i}\mid z_{i^{\prime}}=\beta\right),\forall z_{i},z_{i}^{\prime}\in G,i\neq i^{\prime},\alpha\neq\beta.$
Prompt: The pattern of change is associated with a group. The first two rows of images show the same pattern of change despite other variations in another group. The pattern of change in the last two rows of images should change given other variations. What is the pattern of change? Write in a sentence. If there is no clear pattern, just write “No clear explanation".
Formula:
$p\left(z_{i}\mid p_{k}=\alpha\right)\neq p\left(z_{i}\mid p_{k}=\beta\right),\forall z_{i}\in G_{k},\alpha\neq\beta.$
$p\left(z_{j}\mid p_{k}=\alpha\right)=p\left(z_{j}\mid p_{k}=\beta\right),\forall z_{j}\notin G_{k},\alpha\neq\beta.$
[…]
Formula:
$p\left(z_{i}\mid t=t_{1}\right)\neq p\left(z_{i}\mid t=t_{2}\right),\forall t_{1}\neq t_{2}$.
$p\left(z_{i}\mid\cdot\right)"$ is “pattern of change"""", “$t=t_{1},t=t_{2},\forall t_{1}\neq t_{2}"$ is “different timesteps"""", “"absent"$\neq$"≠ " is “change"""". Please give a prompt in plain language.
Appendix D More Qualitative Evaluation Results
We provide more qualitative evaluation results: (1) the disentanglement bias: DDPM in Figure 5, and $\beta$-TCVAE in Figure 6; (2) the combination bias: CAVAE in Figure 7; (3) the conditional bias: Stable Diffusion in Figure 8 and CSVAE in Figure 9.
Figure 5: Example of generated explanations for DDPM under the disentanglement bias.
Figure 6: Example of generated explanations for $\beta$-TCVAE under the disentanglement bias.
Figure 7: Example of generated explanations for CSVAE under the combination bias.
Figure 8: Example of generated explanations for Stable Diffusion under conditional bias.
Figure 9: Example of generated explanations for CSVAE under conditional bias.
Figure 10: Visualization of the generated explanations with the inductive bias prompt for the combination bias.
Figure 11: Visualization of the generated explanations with the inductive bias prompt for the conditional bias.
Figure 12: Visualization of the generated explanations without the inductive bias prompt for the conditional bias.
Appendix E Visualization of the Generated Explanations
In this section, we provide visualization examples of the explanations generated by our proposed LatentExplainer framework under different inductive biases: combination bias with the inductive bias prompt in Figure 10, conditional bias with the inductive bias prompt in Figure 11 and without the inductive bias prompt Figure 12. Each figure demonstrates the patterns of change in generated image sequences along specific latent directions, accompanied by the corresponding explanations.
Model Method AFHQ LSUN-Church CelebA-HQ
[…]
DDPM (Disentanglement Bias) GPT-4o Baseline 3.08 14.83 14.22 8.78 87.55 -3.39 5.52 23.06 16.85 11.23 88.47 -3.41 0.11 3.54 4.44 1.22 85.97 -3.16
+ LatentExplainer w/o IB 2.14 11.49 12.01 7.82 87.17 -3.41 8.78 27.01 20.15 14.08 88.85 -3.37 0.19 4.17 4.34 1.80 85.88 -3.16
+ LatentExplainer w/o UQ 16.12 31.57 31.74 21.70 89.99 -3.12 21.19 37.03 29.40 23.65 90.49 -3.19 13.11 27.38 25.97 14.03 89.15 -2.92
+ LatentExplainer 25.91 37.84 41.52 29.87 91.48 -2.97 30.92 44.21 36.39 29.23 91.80 -3.06 18.49 35.27 32.76 18.81 90.30 -2.90
Stable Diffusion (Conditional Bias) GPT-4o Baseline 7.61 26.32 21.52 17.19 90.04 -3.32 7.80 26.73 23.40 14.28 89.44 -3.46 5.79 23.89 20.44 12.68 89.63 -3.04
+ LatentExplainer w/o IB 9.12 26.59 22.46 16.28 89.77 -3.29 10.07 28.31 25.28 15.85 89.92 -3.42 5.87 24.06 21.09 12.17 89.15 -3.05
+ LatentExplainer w/o UQ 12.36 27.73 22.64 20.01 90.03 -3.18 17.19 35.33 30.50 21.76 90.73 -3.31 15.01 35.43 27.96 23.99 91.07 -2.87
+ LatentExplainer 13.27 29.98 23.56 20.08 90.17 -3.19 18.33 38.14 32.46 25.01 90.99 -3.23 18.50 40.85 33.90 23.90 91.75 -2.82
Table 5: Quantitative results for diffusion models across datasets. B represents BLEU@4, R represents ROUGE-L, M represents METEOR, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.
Model Method 3DShapes CelebA-HQ dSprites
[…]
$\beta$-TCVAE (Disentanglement Bias) GPT-4o Baseline 5.51 29.77 19.12 7.92 89.18 -3.13 11.09 37.14 28.80 17.24 90.55 -3.07 0.00 16.19 10.38 7.04 87.17 -3.22
+ LatentExplainer w/o IB 5.41 31.66 22.28 10.07 89.44 -3.13 6.14 32.73 25.33 14.04 89.92 -3.11 0.00 18.73 14.42 7.54 87.42 -3.20
+ LatentExplainer w/o UQ 16.99 37.37 30.72 22.89 90.83 -2.80 21.95 48.84 39.70 22.60 92.09 -2.88 17.16 37.19 31.11 22.40 90.18 -2.89
+ LatentExplainer 25.40 49.06 32.72 22.78 91.90 -2.75 29.93 55.68 44.32 30.17 93.55 -2.77 12.55 37.30 29.86 21.69 90.28 -2.87
CSVAE (Combination Bias) GPT-4o Baseline 6.93 25.07 19.04 9.70 89.42 -2.87 10.44 37.42 28.51 15.06 90.53 -3.01 0.00 21.61 13.17 6.99 87.19 -3.17
+ LatentExplainer w/o IB 14.22 28.57 25.12 11.79 90.02 -2.85 11.96 37.02 28.51 17.49 90.50 -3.01 0.00 16.36 10.12 7.26 86.96 -3.24
+ LatentExplainer w/o UQ 34.55 39.62 33.42 30.30 90.83 -2.62 14.08 38.68 30.13 22.49 91.11 -2.87 0.00 28.49 19.19 15.65 89.78 -2.91
+ LatentExplainer 36.18 43.58 36.60 36.75 91.72 -2.52 25.34 45.18 39.82 26.96 92.09 -2.81 16.03 35.85 29.23 21.82 90.05 -2.90
CSVAE (Conditional Bias) GPT-4o Baseline 10.97 30.55 23.90 8.83 90.10 -3.02 8.46 28.78 21.93 15.78 89.62 -2.98 0.00 9.62 5.27 1.03 85.74 -3.34
+ LatentExplainer w/o IB 19.28 39.90 35.47 15.74 90.87 -2.88 8.35 28.75 18.45 12.11 89.60 -2.97 0.00 11.38 9.35 5.26 86.47 -3.29
+ LatentExplainer w/o UQ 16.73 32.28 25.17 16.19 89.84 -2.86 13.35 37.20 25.40 20.05 89.89 -2.92 5.39 19.20 15.65 6.36 86.94 -3.26
+ LatentExplainer 25.71 40.00 31.59 20.67 90.88 -2.79 24.21 42.88 36.19 19.61 91.21 -2.79 21.17 34.99 28.71 20.71 89.49 -3.08
Table 6: Quantitative results for VAE models across datasets. B represents BLEU@4, R represents ROUGE-L, M represents METEOR, S represents SPICE, BS represents BERTScore, and BAS represents BARTScore.